stud
bud
Sign up
Log in
Home
Library
Explore
SAT
More
Probability Vocab Quiz Flashcards | StudBud
Explore
/
Probability Vocab Quiz
Probability Vocab Quiz
13 terms
·
StudBud Administration
·
21 plays
A Quiz of probability in Statistics
Flashcards
13 cards
F
Practice quiz
P
Host live
L
Questions
13
Answers
1
What does General Multiplication Rule require
That A and B are dependent
That A and B are disjoint
That A and B are independent
That both A and B can be made into a Venn Diagram
2
This Symbol:
∩
means...
intersection
union
comparative
Or
3
What does the Addition Rule require?
Disjoint events
Independent Events
That events are equally likely
Sequential Events
4
a compliment is
A mutually exclusive event
The probability that an event will not occur
An event that would make its counterpart impossible
Telling a Walrus you're jealous of his mustache
5
disjoint is ...
when two events are independent
When separate events can happen together
When separate events cannot occur simultaneously
When separate events never effect each others probability
6
Conditional Probability: P(B l A) means ...
The probability of A given B
The probability of B given A
The probability of B and A
The probability of B if A is disjoint
7
P(A) = 1 - P(A
c
) is
Law of Large Numbers
Addition rule
Compliment rule
Multiplication rule
8
independent is ...
When separate events cannot occur simultaneously
Being financial dependent on your parents
When separate events effect each others future outcomes
When separate events do not effect each others probabilities
9
Sample space is:
the set of combined outcomes
theoretical probability
the set of all possible outcomes
An area at Costco used for tasting things
10
The law of large numbers means
Large numbers skew experiment results
The predicted probability of an event if occurs many times
The probability remains constant
Large numbers should not occur in probability answers
11
Probability always adds to:
100
0
10
1
12
Compliment Rule means:
Two intersecting events should be multiplied
Apple (a person) is confusingly named
The probability of A is [ 1 - its compliment ]
Two separate events should be added
13
When does drawing without replacement matter?
Only when you have a very large population
Only when you are looking at one event
Only with smaller groups of experimental units
Only when bias is involved